Calculate the pH at 25°C of 249.0 mL of a buffer solution that is 0.220 M NH4Cl and 0.220 M NH3 before and after the addition of 1.30 mL of 6.0 M HNO3. (The pKa for NH4+ = 9.75)

1) pH before =
2) pH after =
Answer #1

1)
use:
pH = pKa + log {[conjugate base]/[acid]}
= 9.75+ log {0.220/0.220}
= 9.75

Answer: 9.75

2)

mol of HNO3 added = 6.0M *1.3 mL = 7.80 mmol

NH3 will react with H+ to form NH4+

Before Reaction:
mol of NH3 = 0.22 M *249.0 mL
mol of NH3 = 54.78 mmol

mol of NH4+ = 0.22 M *249.0 mL
mol of NH4+ = 54.78 mmol

after reaction,
mol of NH3 = mol present initially - mol added
mmol of NH3 = (54.78 - 7.80) mmol
mol of NH3 = 46.98 mmol

mol of NH4+ = mol present initially + mol added
mol of NH4+ = (54.78 + 7.80) mmol
mol of NH4+ = 62.58 mmol

since volume is both in numerator and denominator, we can use mol instead of concentration
use:
pH = pKa + log {[conjugate base]/[acid]}
= 9.75+ log {46.98/62.58}
= 9.625

Answer: 9.62